OpenStudy (anonymous):

Can someone help me and explain it to me please!!! 4m^2+3m=1 quadratic formula

OpenStudy (abmon98):

4m^2+3m-1=0 is the in the form am^2+bm+c. where a=4, b=3 and c=-1. now use the quadratic formula \[m=-b+-\sqrt{b^2-4ac}/2a\] plugging in a,b and c.

OpenStudy (anonymous):

ok thank u so much

OpenStudy (anonymous):

is that going to be 25?

OpenStudy (danjs):

The -b is also divided by 2a. \[m = \frac{ -b \pm \sqrt{b^2-4*a*c} }{ 2a }\]

OpenStudy (danjs):

Doing a quick factoring... 4m^2 + 3m - 1 = (m+1)*(4m-1) = 0 (m+1) = 0 or (4m-1)=0 The two values for m from the quadratic formula should come out to , m=-1 or m=1/4, if you did the math correctly.